# Beyond Specialization: Assessing the Capabilities of MLLMs in Age and Gender Estimation

### v1

@@ 4.4 Results
Model Test Dataset Age MAPE, % $\downarrow$ Age MAE $\downarrow$ Age CS@5, % $\uparrow$ Gender Acc, % $\uparrow$
MiVOLOv2 IMDB-clean 11.01 3.97 71.16 99.68
384x384 LAGENDA 12.06 3.65 74.48 97.99
@@
Model Age MAPE, % $\downarrow$ Age MAE $\downarrow$ Age CS@5, % $\uparrow$ Gender Acc, % $\uparrow$
ChatGPT4V [29] 19.95 7.07 $\pm$ 6.05 48.08 91.35
LLaVA-NeXT 34B vanilla [23] 22.16 9.23 $\pm$ 8.45 42.31 96.15
MiVOLOv2 384x384 19.82 6.26 $\pm$ 6.17 53.67 96.17
ShareGPT4V 7B fine-tuned 19.36 7.01 $\pm$ 6.67 53.85 95.19
ShareGPT4V 13B fine-tuned 18.27 6.79 $\pm$ 6.42 57.69 97.12
Table 7: Comparison of performance on the Wild104 benchmark: observed standard deviations indicate models instability in average errors.
Figure 4: Failure case. Through manual analysis of ChatGPT’s gender misclassification, we have discovered that it sometimes errs with examples where men have long hair. This is a jokingly artificial illustrative example of such failure cases. Both MiVOLO and the fine-tuned as well as vanilla LLaVA models remain stable.
@@ 5 Conclusions
This study aimed to assess the efficacy of cutting-edge specialized models in comparison to Multimodal Large Language Models (MLLMs) for age and gender estimation tasks.
@@
Our findings reveal a nuanced view. MLLMs, despite not being explicitly trained for facial or bodily analyses to deduce personal attributes, exhibit exceptional capabilities. Notably, models such as ChatGPT stand out by harnessing vast amounts of visual information and training on extensive datasets, demonstrating significant proficiency in tasks beyond their original design. Our research indicates that with minor adjustments, MLLMs can achieve or even surpass the performance of specialized models, suggesting a potential shift towards versatile, general-purpose networks in computer vision. This hints at a future where Multimodal Small Language Models (MSLMs) could potentially replace current specialized models. The flexibility of language models, reducing the need for task-specific models, offers significant advantages for a wide range of applications, especially in scenarios where computational resources and inference speed are not primary concerns. However, in specific data subsets, a well-designed specialized model may still outperform a fine-tuned MLLM. It’s important to note, for the time being, the computational cost of MLLMs cannot be directly compared to that of specialized models — the difference can span thousands of times. Our analysis identified ChatGPT4V as the most precise MLLM for age estimation across numerous benchmarks, despite encountering challenges such as occasional refusal to process images with significant losses ($>21\%$ for our data) and the need for Dynamic High Resolution, which demands a higher computational budget. While this comparison may not be entirely equitable, it highlights the model’s capabilities in ’maximum power mode’. Among open-source alternatives, LLaVA-NeXT 34B leads in this area. The same time, specialized model MiVOLO surpasses all general-purpose MLLMs in age estimation. However, for certain data segments and metrics, fine-tuned specialized versions of LLaVA prove more effective. Such fine-tuned MLLMs present a promising solution for many tasks where computational cost is not a primary concern. Compared to the intensive and expert-driven training required for MiVOLO, fine-tuning an MLLM is considerably simpler, requiring only the same dataset as the specialized model and minimal expertise. Original hyperparameters and losses can be used.
@@
The study highlights the superior performance of MLLMs in gender identification tasks even without any fine-tuning, surpassing that of specialized models. This emphasizes the significance of high-level feature recognition and contextual understanding in this task, where nearly all MLLMs excel. However, ChatGPT-4V stands out for its subpar performance, possibly due to an overemphasis on certain features like hair, which might be influenced by its training data or safety mechanisms. For visualizations, refer to Figure 4. The opaque nature of its development process hampers definitive conclusions. While MLLMs face challenges in maintaining task-specific performance consistency, experiencing notable fluctuations with minor image variations, specialized models show greater stability.
In conclusion, despite current challenges such as high inference costs and other obstacles that may impede their deployment in production environments, the future of technology appears distinctly multimodal. Advances in neural network research and hardware development are progressively making the practical application of multimodal language models more feasible, enhancing their abilities while reducing operational costs. It is expected that the use of versatile and powerful MLMs in production settings will grow, steadily reducing the dependence on specialized models for a wide range of tasks.

### v4

@@ 4.4 Results
Model Test Dataset Age Gender
MAPE, % $\downarrow$ MAE $\downarrow$ CS@5, % $\uparrow$ Acc, % $\uparrow$
$MiVOLO_{384}$ IMDB-clean 11.01 3.97 71.16 99.68
LAGENDA 12.06 3.65 74.48 97.99
@@
Model Age Gender
MAPE, % $\downarrow$ MAE $\downarrow$ CS@5, % $\uparrow$ Acc, % $\uparrow$
ChatGPT4V [30] 19.95 7.07 48.08 91.35
ChatGPT4O [30] 16.11 6.07 58.65 91.35
LLaVA-NeXT 34B vanilla [24] 22.16 9.23 42.31 96.15
$MiVOLO_{384}$ 19.82 6.26 53.67 96.17
ShareGPT4V 7B fine-tuned 19.36 7.01 53.85 95.19
ShareGPT4V 13B fine-tuned 18.27 6.79 57.69 97.12
Table 10: Comparison of performance on the Wild104 benchmark.
Figure 4: Failure case: Through manual analysis of ChatGPT’s gender misclassification, we have discovered that it sometimes makes mistakes with examples where men have long hair. This is a humorously artificial illustrative example of such failure cases. Both MiVOLO and the fine-tuned as well as vanilla LLaVA models remain stable.
@@ 5 Conclusions
This study aimed to assess the efficacy of cutting-edge specialized models in comparison to MLLMs for age and gender estimation tasks.
@@
Our findings reveal a nuanced view. MLLMs, despite not being explicitly trained for facial or bodily analyses to deduce personal attributes, exhibit exceptional capabilities. Notably, models such as ChatGPT stand out by harnessing vast amounts of visual information and training on extensive datasets, demonstrating significant proficiency in tasks beyond their original design. Our analysis identified ChatGPT-4O as the most precise MLLM for age estimation across numerous benchmarks, despite encountering challenges such as occasional refusal to process images with significant losses ($>21\%$ for our data) and the need for Dynamic High Resolution, which demands a much higher budget. While this comparison may not be entirely equitable, it highlights the model’s capabilities in ’maximum power mode’. Among open-source alternatives, LLaVA-NeXT 34B leads in this area. At the same time, the improved specialized model $MiVOLO_{384}$ surpasses all general-purpose open-source MLLMs in age estimation. However, for certain data segments and metrics, fine-tuned specialized versions of LLaVA prove more effective. Such fine-tuned MLLMs present a promising solution for many tasks where computational cost is not a primary concern. Compared to the tricky and expert-driven training required for MiVOLO, fine-tuning an MLLM is considerably simpler, requiring only the same dataset as the specialized model and minimal expertise. Original hyperparameters and losses can be used.
@@
The study highlights the superior performance of MLLMs also in gender identification tasks even without any fine-tuning, surpassing that of specialized models. This emphasizes the significance of high-level feature recognition and contextual understanding in this task, where nearly all MLLMs excel. However, ChatGPT-4V and even the newest ChatGPT-4O stand out for their subpar performance, possibly due to an overemphasis on certain features like hair, which might be influenced by its training data or safety mechanisms. For visualizations, refer to Figure 4. The opaque nature of its development process hampers definitive conclusions.
Overall, our research indicates that with minor adjustments, open-source MLLMs can achieve or even surpass the performance of specialized models, suggesting a potential shift towards versatile, general-purpose networks in computer vision. The flexibility of language models offers significant advantages for a wide range of applications, especially in scenarios where computational resources and inference speed are not primary concerns. However, it is important to note that, for the time being, the computational cost of MLLMs cannot be directly compared to that of specialized models — the difference can span thousands of times. Possibly, in the future, Multimodal Tiny Language Models could turn the tables.